# TaSL: Task Skill Localization and Consolidation for Language Model Continual Learning

### v1

@@ III-B Low-Rank Adaption
Low-Rank Adaption (LoRA) [7] is a parameter-efficient fine-tuning method to adapt LLMs to novel tasks. It assumes that the changes of parameters lie in a low-rank space when the model is fully fine-tuned on a downstream task. This approach eliminates the necessity of fine-tuning the entire model, thereby improving computational efficiency and resource utilization.
Inspired by the low-rank internal dimensionality [41], LoRA hypothesizes the updates to the weights also has a low “intrinsic rank” during adaptation. For a pre-trained weight matrix $W^{(0)}\in\mathbb{R}^{out\times in}$ that takes $x$ as input, LoRA modifies the output $h$ of $W^{(0)}$ with a low-rank decomposition:
@@
while $B\in\mathbb{R}^{out\times r}$, $A\in\mathbb{R}^{r\times in}$, and the rank $r\ll\min(in,out)$. $A$ is initialized from a random Gaussian distribution and $B$ is initialized with all zeros. During the training stage, $W^{(0)}$ is fixed, and only $BA$ is updated. Due to its superior performance, LoRA has become one of the most popular PEFT methods in NLP community.
@@ IV-A Overview
TaSL includes two key components: (i) Skill Localization, utilizing a group-wise importance metric to accurately identify the importance distribution of parameters across tasks, and (ii) Skill Consolidation, which employs a novel fine-grained model averaging strategy to integrate model weights from both current and past tasks for effective knowledge transfer.
Figure 2 provides a comprehensive overview of our proposed TaSL framework, with the following subsections detailing each component and the corresponding extensions.
Figure 2: Left: Illustration of reconstructing LoRA as fine-grained skill units. Right: Overview of TaSL. Step 1: We compute the importance scores of skill units for the current task $k$ using our importance-aware skill localization method during fine-tuning. Step 2: Based on a categorical model averaging strategy, the skill consolidation mechanism merges the model $\hat{f}_{k-1}$, which accumulates knowledge of all previous tasks, with the current task’s model $f_{k}$. The integration is guided by the importance distributions of skill units across various tasks. This process can be repeated with the introduction of each subsequent task.
@@ IV-B Fine-Grained Skill Unit
Before we localize the importance of parameters, we need to define a structure that helps us better organize the knowledge stored in the model. This structure will provide a clearer understanding of how to mitigate CF and achieve KT. Therefore, we refer to the basic unit of stored skills or knowledge in the model as a “skill unit.” Based on different usage scenarios, we offer the following two structures for skill units:
@@ IV-B1 Matrix-Level Skill Unit
In this division strategy, we define skill units as individual matrices in the model, such as the query or key matrices in the self-attention layer or the A matrix in LoRA. This approach aligns with the original TaSL framework. The advantage of this division is that it is model-agnostic, making it applicable to both traditional full-parameter fine-tuning and the latest parameter-efficient fine-tuning methods, such as LoRA. This provides a higher degree of generalizability.
@@ IV-B2 LoRA-Tailored Skill Unit
While the matrix-level division is simple and intuitive, it did not sufficiently consider the intra-matrix redundancy and inter-matrix dependencies. To address these limitations and the inefficiencies noted in recent work [19], which treats each LoRA adapter as a container for task-specific knowledge, we decompose the model matrices into new, finer-grained “skill units” based on parameter dependencies tailored for LoRA. The construction process for these LoRA-tailored skill units is as follows.
@@
where skill unit $u_{i}$ is a matrix obtained by the product of the vectors $b_{i},a_{i}$. Therefore, LoRA can be viewed as a fusion of knowledge learned from multiple skill units:
@@
This division strategy effectively decomposes the parameters of a single LoRA adapter, reducing the redundancy in the traditional approach of employing separate LoRA adapters for each task [19]. To ensure that different skill units within the same layer learn distinct latent semantic features, we apply a regularization term as introduced by [39]:
@@
where $I$ is an identity matrix, this forces $A$ and $B$ to be orthogonal after training.
@@ IV-C Importance-aware Skill Localization
To calculate the importance of each skill unit $u$, we introduce a group-wise metric that also addresses the significant computational and storage burdens associated with previous parameter-level importance calculation methods [42]:
@@
where $w_{ij}$ denotes the trainable parameters, and $in\times out$ represents the total parameter count in a skill unit $u$. $\mathcal{I}(u)$ measures the collective importance of all parameters within each skill unit, where higher values signify increased importance. The importance function $s(\cdot)$ for individual parameters, inspired by the pruning community [43], is quantified by measuring the impact of its removal on the loss. Specifically, to estimate the importance of $W_{i}$, we calculate the change in loss when the parameter is zeroed out using:
@@
where $H$ is the hessian matrix, $W_{i}$ represents the $i$-th parameter in $W$, and $\mathcal{L}$ represents the next-token prediction loss. If removing a parameter has a significant influence, then the model is sensitive to it, and we should retain it [44]. Based on Eq. (7), we can derive the following two metrics for calculating parameter importance:
@@ IV-C1 First-Order Gradient-Based Metric
In original TaSL framework, we followed the standard approach, defining $s(\cdot)$ as the magnitude of the gradient-weight product:
@@
This metric only considers the first-order gradient term in the importance calculation. However, the importance of a parameter is also reflected in the second-order gradients (Hessian matrix). Ignoring second-order gradient information may lead to biases in localization, thus affecting model performance. Therefore, we propose the following new metric.
@@ IV-C2 Second-Order Gradient Approximation Metric
Direct computation of the Hessian matrix for LLMs is impractical due to its $\mathcal{O}(N^{2})$ complexity. To address this, we propose a second-order gradient approximation for the importance metric, which balances the efficiency of first-order methods with the precision of second-order gradients. To reduce the computational complexity, the diagonal of the Hessian $H_{ii}$ can be approximated by the Fisher information matrix [45], and the importance can be defined as:
@@
However, calculating the importance as specified in Eq. (9) across the entire training set introduces significant challenges, as model training typically only has access to mini-batch data. This means the metric is subject to variability due to stochastic sampling and training dynamics, introducing large uncertainty in estimating sensitivity. To mitigate this, we propose a more reliably importance metric based on sensitivity smoothing and uncertainty quantification [39]:
@@
where $\alpha_{1}$ and $\alpha_{2}$ are smoothing factors, and $t$ is the iteration number. $\bar{I}^{(t)}$ represents smoothed sensitivity and $\bar{U}^{(t)}$ is the uncertainty term quantified by the local variation between $I^{(t)}$ and $\bar{I}^{(t)}$. Using the exponential moving average of the importance metric, we can retain and explore the trajectory gradient for a longer time, providing a more robust and precise importance assessment. Importance is then determined by:
@@
To compute the importance score of each skill unit for current task $\mathcal{T}_{k}$, we employ Eq. (6) during fine-tuning. The model $f$ with $n$ skill units is denoted as $\mathcal{U}=\left\{u_{1},\ldots,u_{n}\right\}$, with their importance scores for task $\mathcal{T}_{k}$ denoted by $\mathcal{I}(\mathcal{U}_{k})\in\mathbb{R}^{n}$. The detailed computation process is provided in Algorithm 1.
@@
After computing importance scores for each skill unit at the current task $\mathcal{T}_{k}$, it is essential to compare these with scores from all previously learned tasks to distinguish between task-specific and task-shared skill units. To avoid the inefficiency of storing scores for each past task, we aggregate importance scores from all prior tasks into a cumulative score for tasks up to $\mathcal{T}_{k-1}$. This method allows for the iterative refinement of accumulated scores without separately saving past task scores. The skill units with these cumulative scores up to $\mathcal{T}_{k-1}$ are denoted as $\hat{\mathcal{U}}_{k-1}$, calculated using:
@@
where $\beta\in[0,1]$, and Norm($\cdot$) normalizes importance scores to the [0, 1] range, thus resolving discrepancies across models. The initial scores, $\mathcal{I}(\hat{\mathcal{U}}_{1})$, are set to be equal to $\mathcal{I}(\mathcal{U}_{1})$. Following this, the importance distribution for skill units up to task $\mathcal{T}_{k-1}$ is combined with that of the current task, $\mathcal{T}_{k}$, to facilitate the skill consolidation process.
@@ IV-D Skill Consolidation
After skill localization, the subsequent vital phase involves consolidating the knowledge in the skill unit into a unified framework. This process demands a sophisticated model averaging approach considering various factors to optimize task performance. Traditional coarse-grained model averaging assumes that all model weights are equally important for the training task [46, 47], which can be written as the following iterative computation format:
@@
However, coarse-grained methods may overemphasize weights that are irrelevant to the current task, contaminating previously acquired task-specific knowledge and leading to forgetting. To address this issue, we introduce two fine-grained averaging strategies that focus on skill units rather than the entire model. Our approach distinguishes between task-shared and task-specific skill units, applying weighted averaging to parameters within each unit. This tailored method provides a customized solution to meet the distinct needs of each task.
@@ IV-D1 Static Weighted Consolidation
This strategy, used in the original TaSL framework, begins by setting importance thresholds $\delta$ using quantiles to select the top 20% of skill units based on importance scores. A skill unit $u_{i}^{k}$ is deemed important (denoted as $(u_{i}^{k})^{+}$) if its score $\mathcal{I}(u_{i}^{k})$ is above $\delta_{k}$, and unimportant ($(u_{i}^{k})^{-}$) otherwise.
Our static weighted consolidation strategy customizes parameter combination for each skill unit, based on its importance under different tasks, as follows:
@@
This approach performs the element-wise adjustment of parameters within each skill unit based on its relevance to previous and current tasks, using hyperparameter $\gamma$ to control their influences.
@@ IV-D2 Adaptive Weighted Consolidation
Building upon the original TaSL framework, we propose an Adaptive Weighted Consolidation strategy as our latest extension. In contrast to the manual setting of multiple hyperparameters required in the original method, this adaptive strategy simplifies the process by automatically adjusting to various scenarios, making the averaging process more efficient and flexible. In detail, for a specific skill unit $u_{i}$, the weighting coefficients are:
@@
By using Eq. (15) or Eq. (17) in our skill consolidation phrase, it effectively addresses the challenges in CL:
  • If $\mathcal{I}(\hat{u}^{k-1}_{i})$ is high while $\mathcal{I}(u^{k}_{i})$ is low, it indicates that the skill unit $u_{i}$ is more important for the history task. According to Eq. (15) or (17), we maintain the previous task-specific knowledge stored in $\hat{u}^{k-1}_{i}$ untouched to prevent the contamination of historical knowledge with task-irrelevant information in $u^{k}_{i}$, thereby mitigating CF.
  • If $\mathcal{I}(\hat{u}^{k-1}_{i})$ is low and $\mathcal{I}(u^{k}_{i})$ is high, it suggests that the skill unit $u_{i}$ is more important for the current task. Since the model’s training initializes from the endpoint of the previous task, the historically learned knowledge is utilized to enhance performance on the current task. So we ensure the integrity of the parameters in $u^{k}_{i}$, which stores current task-specific knowledge, thereby facilitating forward KT.
  • If both $\mathcal{I}(\hat{u}^{k-1}_{i})$ and $\mathcal{I}(u^{k}_{i})$ are high, it indicates that the skill unit $u_{i}$ is important for both the previous and current tasks. Consequently, we integrate newly acquired knowledge into this task-shared skill unit to enable backward KT.
  • If both $\mathcal{I}(\hat{u}^{k-1}_{i})$ and $\mathcal{I}(u^{k}_{i})$ are low, it indicates the skill unit is not significantly relevant to any task. In this case, simply averaging the parameters within this unit is sufficient.
Skill consolidation is performed before starting a new task in CL, utilizing the averaged model for subsequent task initialization. Only the importance scores of $\hat{\mathcal{U}}_{k-1}$ and $\mathcal{U}_{k}$ are retained for use between tasks, starting with $\hat{\mathcal{U}}_{1}=\mathcal{U}_{1}$ estimated from $f_{1}$ on $D_{1}$. Detailed implementation of entire TaSL algorithm is provided in Algorithm 2.
@@ V TaSL with Memory Replay: TaSL-M
To adapt TaSL to scenarios where historical data can be used, we propose a memory replay-enhanced version, TaSL-M, to further improve model performance. The implementation process is shown in Figure 3.
Figure 3: Overview of TaSL-M.
After training on task $\mathcal{T}_{k}$ using the skill localization and skill consolidation, we obtain the model $\hat{f}_{k}$. Before the next task arrives, we introduce a replay stage, where the model $\hat{f}_{k}$ is fine-tuned using historical data stored in the memory buffer $\mathcal{M}_{<k}$. This process results in a new model $\hat{f}^{m}_{k}$, which helps recover forgotten knowledge. Specifically, during the replay stage, we do not fine-tune the parameters of all skill units using historical data, as this could negatively impact the performance on the current task. Instead, we fix the current task-specific skill units and fine-tune only the remaining parameters. This approach achieves a better balance between retaining previous knowledge and excelling in new tasks.

### v2

@@ III-B Low-Rank Adaption
Low-Rank Adaption (LoRA) [7] is a parameter-efficient fine-tuning method to adapt LLMs to novel tasks. It operates under the assumption that parameter changes during full fine-tuning on a downstream task reside within a low-rank space. This approach allows for not having to fine-tune the entire model, thus significantly improving computational efficiency and resource utilization. Inspired by the low-rank internal dimensionality [34], LoRA hypothesizes the updates to the weights also has a low “intrinsic rank” during adaptation. For instance, consider a pre-trained weight matrix $W^{(0)}\in\mathbb{R}^{out\times in}$ that takes $x$ as input, LoRA modifies the output $h$ of $W^{(0)}$ with a low-rank decomposition:
@@
while $B\in\mathbb{R}^{out\times r}$, $A\in\mathbb{R}^{r\times in}$, and the rank $r\ll\min(in,out)$. $A$ is initialized from a random Gaussian distribution and $B$ is initialized with all zeros. During training, $W^{(0)}$ remains fixed, and only $BA$ is updated. Due to its effectiveness, LoRA has emerged as one of the most favored PEFT methods within the NLP community [35].
@@ IV-A Overview
TaSL includes two key components: (i) Skill Localization, utilizing a group-wise importance metric to precisely identify the importance distribution of parameters across tasks, and (ii) Skill Consolidation, which employs an innovative fine-grained model averaging strategy to integrate model weights from both current and previous tasks for effective knowledge transfer.
Figure 2 presents a detailed overview of our proposed TaSL framework, with the following subsections elaborating on each component and the corresponding extensions.
Figure 2: Left: Depiction of reconstructing LoRA as fine-grained skill units. Right: Overview of TaSL. Step 1: We compute the importance scores of skill units for the current task $k$ using our importance-aware skill localization method during fine-tuning. Step 2: Based on a categorical model averaging strategy, the skill consolidation mechanism merges the model $\hat{f}_{k-1}$, which accumulates knowledge of all previous tasks, with the current task’s model $f_{k}$. This integration is strategically guided by the importance distributions of skill units across different tasks. This iterative process continues with the addition of each new task.
@@ IV-B Fine-Grained Skill Unit
Before localizing the importance of parameters, it is essential to define a structure that facilitates better organization and understanding of the knowledge stored within the model. This structure aids in addressing CF and enhancing KT. We define this basic structural element of stored skills or knowledge in the model as a “skill unit.” Depending on different usage scenarios, we propose two structures for skill units:
@@ IV-B1 Matrix-Level Skill Unit
In this division strategy, we define skill units as individual matrices in the model, such as the query or key matrices in the self-attention layer or the A matrix in LoRA. This approach aligns with the original TaSL framework and offers the advantage of being model-agnostic. It is suitable for both traditional full-parameter fine-tuning and modern parameter-efficient fine-tuning methods like LoRA, ensuring broad generalizability.
@@ IV-B2 LoRA-Tailored Skill Unit
While the matrix-level division is straightforward and intuitive, it does not adequately address intra-matrix redundancy or inter-matrix dependencies.
To address these limitations and the inefficiencies noted in recent work [18], which treats each LoRA adapter as a container for task-specific knowledge, we propose a more refined decomposition of model matrices into new, finer-grained “skill units” based on parameter dependencies specifically tailored for LoRA. The construction process for these LoRA-tailored skill units unfolds as follows.
@@
where each skill unit $u_{i}$ is a matrix formed by the product of the vectors $b_{i},a_{i}$. Thus, LoRA can be conceptualized as an aggregation of knowledge across multiple skill units:
@@
This division strategy significantly reduces the redundancy inherent in the traditional approach of utilizing separate LoRA adapters for each task [18]. To ensure that different skill units within the same layer learn distinct latent semantic features, we incorporate a regularization term as introduced in [36]:
@@
where $I$ is the identity matrix, this forces $A$ and $B$ to be orthogonal after training.
@@ IV-C Importance-aware Skill Localization
To calculate the importance of each skill unit $u$, we introduce a group-wise metric that mitigates the significant computational and storage burdens associated with previous parameter-level importance calculation methods [37]:
@@
where $w_{ij}$ denotes the trainable parameters, and $in\times out$ represents the total parameter count in a skill unit $u$. The function $\mathcal{I}(u)$ measures the collective importance of all parameters within each skill unit, with higher values indicating greater importance. The importance function $s(\cdot)$ for individual parameters, inspired by the pruning community [38], is quantified by measuring the impact of its removal on the loss. Specifically, to estimate the importance of $W_{i}$, we measure the change in loss when the parameter is zeroed out using:
@@
where $H$ is the Hessian matrix, $W_{i}$ is the $i$-th parameter in $W$, and $\mathcal{L}$ is the next-token prediction loss. If removing a parameter has a significant influence, then the model is sensitive to it, and we should retain it. Based on Eq. (7), we can derive the following two metrics for calculating importance:
@@ IV-C1 First-Order Gradient-Based Metric
In the original TaSL framework, we defined the importance function $s(\cdot)$ as the magnitude of the gradient-weight product:
@@
This metric primarily considers the first-order gradient term in determining parameter importance. However, the significance of a parameter is also influenced by second-order gradients, represented by the Hessian matrix. Overlooking this second-order gradient information can introduce biases in the localization process, potentially degrading model performance. Therefore, we propose the following new metric.
@@ IV-C2 Second-Order Gradient Approximation Metric
The direct computation of the Hessian matrix for LLMs is impractical due to its $\mathcal{O}(N^{2})$ complexity. To address this, we propose a second-order gradient approximation metric that balances the efficiency of first-order methods with the precision of second-order gradients. To reduce computational demands, we approximate the diagonal of the Hessian, $H_{ii}$, using the Fisher information matrix [39]. The importance is then defined as:
@@
However, calculating the importance as specified in Eq. (9) across the entire training set introduces significant challenges, as model training typically accesses only mini-batch data. This limitation subjects the metric to variability due to stochastic sampling and training dynamics, introducing substantial uncertainty in estimating parameter sensitivity. To mitigate this, we propose a more reliable importance metric based on sensitivity smoothing and uncertainty quantification [36]:
@@
where $\alpha_{1}$ and $\alpha_{2}$ are smoothing factors, and $t$ is the iteration number. $\bar{I}^{(t)}$ represents smoothed sensitivity and $\bar{U}^{(t)}$ is the uncertainty term quantified by the local variation between $I^{(t)}$ and $\bar{I}^{(t)}$. Using the exponential moving average of the importance metric, we can retain and examine the trajectory gradient for a longer time, yielding a more precise importance assessment. Importance is ultimately determined by:
@@
To calculate the importance score of each skill unit for the current task $\mathcal{T}_{k}$, we use Eq. (6) during fine-tuning. The model $f$ with $n$ skill units is denoted as $\mathcal{U}=\left\{u_{1},\ldots,u_{n}\right\}$, with their importance scores for task $\mathcal{T}_{k}$ denoted by $\mathcal{I}(\mathcal{U}_{k})\in\mathbb{R}^{n}$. The detailed computation process is provided in Algorithm 1.
@@
After computing the importance scores for each skill unit at the current task $\mathcal{T}_{k}$, it is crucial to compare these with scores from all previously learned tasks to distinguish between task-specific and task-shared skill units. To streamline this process and avoid the inefficiencies associated with storing scores for each past task, we aggregate importance scores from all prior tasks into a cumulative score for tasks up to $\mathcal{T}_{k-1}$. This approach allows for iterative refinement of accumulated scores without the need to separately store past task scores. The skill units with these cumulative scores up to $\mathcal{T}_{k-1}$ are denoted as $\hat{\mathcal{U}}_{k-1}$, calculated using:
@@
where $\beta\in[0,1]$ and Norm($\cdot$) normalizes the importance scores to the [0, 1] range, ensuring consistency across models. The initial scores, $\mathcal{I}(\hat{\mathcal{U}}_{1})$, are set equal to $\mathcal{I}(\mathcal{U}_{1})$. Following this, the importance distribution for skill units up to task $\mathcal{T}_{k-1}$ is combined with that of the current task, $\mathcal{T}_{k}$, to facilitate the skill consolidation process.
@@ IV-D Skill Consolidation
Following skill localization, the next critical phase is skill consolidation, where knowledge within each skill unit is integrated into a cohesive framework. This process utilizes a sophisticated model averaging approach that accounts for various factors to optimize task performance.
Traditional coarse-grained model averaging operates under the assumption that all model weights hold equal importance for the training task [23, 40], typically implemented through the following iterative computation:
@@
However, coarse-grained methods may overemphasize weights that are irrelevant to the current task, contaminating previously acquired task-specific knowledge and leading to forgetting. To overcome these limitations, we introduce two fine-grained averaging strategies that focus on skill units instead of the entire model. Our method distinguishes between task-shared and task-specific skill units, applying a weighted averaging technique to the parameters within each unit. This refined approach offers a tailored solution, addressing the unique requirements of each task more effectively.
@@ IV-D1 Static Weighted Consolidation
In the original TaSL framework, this consolidation strategy initiates by establishing importance thresholds $\delta$ through quantiles, selecting the top 20% of skill units based on importance scores. A skill unit $u_{i}^{k}$ is considered important (denoted as $(u_{i}^{k})^{+}$) if its score $\mathcal{I}(u_{i}^{k})$ exceeds $\delta_{k}$, and unimportant ($(u_{i}^{k})^{-}$) otherwise.
The static weighted consolidation strategy tailors parameter integration for each skill unit, according to its importance across different tasks. The method is defined as follows:
@@
This approach performs the element-wise adjustment of parameters within each skill unit based on its relevance to previous and current tasks. The hyperparameter $\gamma$ is employed to control their influences.
@@ IV-D2 Adaptive Weighted Consolidation
As a sophisticated extension to the static method, which necessitated the manual setting of multiple hyperparameters, this adaptive strategy simplifies the process by automatically adjusting to various scenarios. This enhancement increases the efficiency and flexibility of the averaging process. Specifically, for a specific skill unit $u_{i}$, the weighting coefficients are determined by:
@@
By applying Eq. (15) or Eq. (17) in our skill consolidation phrase, it effectively addresses the challenges in CL:
  • Mitigating CF: When $\mathcal{I}(\hat{u}^{k-1}_{i})$ is high and $\mathcal{I}(u^{k}_{i})$ is low, this indicates that the skill unit $u_{i}$ is more critical for previous tasks. According to Eq. (15) or (17), we maintain the previous task-specific knowledge in $\hat{u}^{k-1}_{i}$ untouched, preventing contamination from the less relevant $u^{k}_{i}$.
  • Facilitating Forward KT: Conversely, if $\mathcal{I}(\hat{u}^{k-1}_{i})$ is low and $\mathcal{I}(u^{k}_{i})$ is high, it suggests that $u_{i}$ holds greater importance for the current task. As training initiates from the endpoint of the previous task, maintaining the integrity of parameters in $u^{k}_{i}$ leverages historically learned knowledge to enhance performance on the current task.
  • Enabling Backward KT: When both $\mathcal{I}(\hat{u}^{k-1}_{i})$ and $\mathcal{I}(u^{k}_{i})$ are high, $u_{i}$ is vital for both previous and current tasks. Here, we integrate newly acquired knowledge into this task-shared skill unit to support backward KT.
  • Handling Low Relevance: If both $\mathcal{I}(\hat{u}^{k-1}{i})$ and $\mathcal{I}(u^{k}{i})$ are low, indicating minimal relevance to any task, a simple averaging of parameters within this unit is adequate.
Skill consolidation is conducted before starting a new task in CL, utilizing the averaged model for subsequent task initialization. Only the importance scores of $\hat{\mathcal{U}}_{k-1}$ and $\mathcal{U}_{k}$ are preserved for use between tasks, starting with $\hat{\mathcal{U}}_{1}=\mathcal{U}_{1}$ estimated from $f_{1}$ on $D_{1}$. Detailed implementation of entire TaSL algorithm is provided in Algorithm 2.
@@ V TaSL with Memory Replay: TaSL-M
To adapt TaSL for scenarios where historical data are accessible, we introduce a memory replay-enhanced version, TaSL-M, designed to further improve model performance. The implementation process is shown in Figure 3.
Figure 3: Overview of TaSL-M. By fixing the current task-specific skill unit during the replay phase, we can further enhance the performance of TaSL.
Following the training on task $\mathcal{T}_{k}$ using the skill localization and skill consolidation, we obtain the model $\hat{f}_{k}$. Before the next task arrives, a replay stage is introduced, where the model $\hat{f}_{k}$ is fine-tuned using historical data stored in the memory buffer $\mathcal{M}_{<k}$. This process results in a new model $\hat{f}^{m}_{k}$, which helps recover forgotten knowledge.
Specifically, during the replay stage, we avoid fine-tuning the parameters of all skill units with historical data, as this could potentially degrade the model’s performance on the current task. Instead, we maintain the integrity of the current task-specific skill units and selectively fine-tune the remaining parameters. This approach achieves a better balance between retaining previous knowledge and excelling in new tasks.